OpenStudy (anonymous):

Ms.Blackmon invested $8000 part 4% and the rest at 5% per year how much did she invest at each rate if her total return per year on the investment was $380 ?

hero (hero):

x + y = 8000 .04x + .05y = 380

OpenStudy (anonymous):

okay what we do next

hero (hero):

There are a few things you could do, but it might be easier to multiply the first equation by .04: .04x + .04y = 320 .04x + .05y = 380 .04x = .04x 320 - .04y = 380 - .05y .05y - .04y = 380 - 320 .01y = 60 y = 60/.01 y = 6000 x = 2000

OpenStudy (anonymous):

how did u get that

hero (hero):

I just showed you how I got it. Study the steps

OpenStudy (anonymous):

lol my bad